How many liters of water needs to be added to 7.00 g of Magnesium Cyanide to make a 0.500 solution of Magnesium Cyanide?
1
Expert's answer
2020-03-30T10:53:02-0400

At first let's find the amount of Magnesium Cyanide:

"\nu = m/M(Mg(CN)_2) = 7 (g)/76 (g/mol) = 0.092 (mol)"


Then find the volume of water:

"V = \nu/C = 0.092 (mol) /0.500 (mol/L) = 0.184 (L)"


Answer: 0.184 L of water needs to be added to 7.00 g of Magnesium Cyanide to make a 0.500M solution of Magnesium Cyanide